OpenStudy (anonymous):

a class consists of 15 boys and 12 girls. On tuesday , all the boys were present and the number of girls who were represented 40 percent of the class attendance for that day. How many girls were ABSENT on tuesday?

OpenStudy (kc_kennylau):

I don't get the question

OpenStudy (anonymous):

let g = number of girls abent, then 12 - g = 0.4 (15 + 12 - g) g = 2 answer: 2 girls absent on Tuesday